OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to the line y equals three-fifths x minus 1 and passes through the point (-9, 4). A. five-thirds x plus 19 B. three-fifths x plus forty-seven-fifths C. negative five-thirds x minus 11 D. negative five-thirds x plus 19

OpenStudy (anonymous):

\[y=\frac{3}{5}x-1\] though (-9,4)??

OpenStudy (anonymous):

slope will be \[-\frac{5}{3}\] then point slope formula \[y-y_1=m(x-x_1)\]with \[x_1=-9,y_1=4, m=-\frac{5}{3}\]

OpenStudy (anonymous):

-5/3x-11 or -5/3x+19 ?

OpenStudy (anonymous):

first one